OpenStudy (anonymous):

sin 20 .sin 40 .sin80

OpenStudy (raden):

use this formula : sin(a) * sin(60-a) * sin(60+a) = sin(3a) / 4 set a=20, giving us : sin 20 sin 40 sin80 = sin60/4 = (1/2 sqrt(3)) /4 = 1/8 sqrt(3)